# Can Static Black Holes in Massive Gravity Serve as Candidates for Aschenbach-Like Phenomena?

### v1

@@ Abstract
The Aschenbach effect is widely regarded as a manifestation of two quintessential relativistic features: frame dragging and extreme spacetime curvature. Traditionally associated with rotating geometries, this non-monotonic behavior in orbital angular velocity challenges Newtonian intuition. In our previous work, however, we demonstrated that this velocity irregularity is not exclusive to spinning spacetimes. Specifically, we showed that the presence of a stable minimum in the gravitational potential, corresponding to a stable photon sphere, can reproduce Aschenbach-like behavior in static black holes as well. This observation suggests that, even in the absence of rotational frame dragging, curvature alone—if encoded through appropriate geometric extrema—may be sufficient to induce non-monotonic velocity profiles.
@@
Aschenbach effect, photon spheres, Time-like Circular Orbits, black holes, Massive Gravity
I Introduction
The Aschenbach effect, represents a compelling case of deep interplay between fundamental physics and observational astrophysics. This phenomenon is intricately tied to black hole physics, and since black holes constitute a tangible manifestation of general relativity in the physical universe, general relativity serves as the foundational framework for explaining the Aschenbach effect. The theory provides a robust mathematical structure for describing how matter and energy influence spacetime curvature within this context. In other words, The Aschenbach phenomenon is a direct prediction of general relativity in the strong-field regime, which observing it would confirm that spacetime near black holes behaves as Einstein’s equations predict.
@@ I Introduction
The origin of the Aschenbach effect can be traced back to pioneering investigations by researchers exploring relativistic particle motion around rotating (Kerr) black holes1 . While the concept of frame dragging—a hallmark prediction of general relativity—had already been established 2 ; 3 , the systematic analysis of orbital stability near Kerr black holes gained momentum in the late 20th century. Key figures such as Brandon Carter, and later Marek A. Abramowicz, made substantial contributions to the mathematical formalism underpinning these orbital dynamics4 ; 5 ; 6 . However, in 2004, it was Aschenbach who, through numerical simulations, discovered an anomalous drop in orbital velocity for test particles extremely close to the event horizon just prior to plunge7 . This non-monotonic behavior challenges the assumptions of traditional Keplerian flow models, often employed in simplified accretion disk theories.
The emergence of this anomalous oscillatory motion near the innermost stable circular orbit (ISCO) is notably distinct from quasi-periodic oscillations (QPOs). Unlike QPOs—which typically require magnetohydrodynamic (MHD) simulations for accretion flow modeling 8 —the Aschenbach effect is purely kinematic and appears to arise directly from Kerr geometry. Rigorous inferences from Einstein’s field equations demonstrate that the phenomenon originates from the interplay between gravitational forces and rotational effects encoded in the Kerr metric. The Kerr metric, characterized by its angular momentum parameter, governs fundamental properties of rotating black holes, including ergosphere dynamics and frame-dragging. Concurrently, accretion disks around rotating black holes are regulated by complex interactions among gravitational forces, frame-dragging effects, and viscous dissipation processes 9 . Contemporary studies employing Kerr-metric-governed numerical simulations of accretion disks have successfully reproduced key aspects of the Aschenbach effect, thereby validating its theoretical underpinnings10 . Additionally, we know that X-ray emissions from regions proximal to spining black holes provide crucial insights into ISCO properties11 , which are directly influenced by the spin parameter of the central object. Accordingly, based on the analysis of observations conducted using instruments such as NASA’s Chandra X-ray Observatory and ESA’S XMM-Newton satellite, one can be very hopeful of obtaining spectral signatures consistent with predictions arising from the Aschenbach effect 12 .
In light of the preceding discussion, it is timely to address the core question:
What, precisely, is the Aschenbach effect, and what fundamentally constitutes this effect? To answer this rigorously, we must introduce and examine the fundamental physical elements that underpin the phenomenon:
$\clubsuit$ The ISCO :
represents the smallest radius at which a test particle can maintain a stable circular orbit around a black hole. Beyond this threshold, orbital trajectories become unstable — minor perturbations (such as a small loss of angular momentum) trigger inward plunges.
@@
ISCO effectively delineates the inner edge of accretion disks, Matter interior to this radius cannot sustain circular motion and undergoes rapid radial infall. In rotating (Kerr) spacetimes, ISCO properties depend on orbital orientation:
Prograde Orbits (co-rotating with spin) and Retrograde Orbits (counter-rotating)13 .
$\clubsuit$ Frame Dragging:
@@
The rotation of a Kerr black hole ”drags” the surrounding spacetime, altering the dynamics of nearby matter. This relativistic effect is strongest near the event horizon and weakens with distance. It is quantified by the spin parameter $a=J/M$, where $J$ is the angular momentum and $M$ is the mass. For maximal spin ($a/M=1$), frame dragging reaches its extremum; for non-spinning Schwarzschild black holes ($a=0$), the effect vanishes. Prograde orbits gain angular momentum through frame dragging, modifying expected kinematics.13
$\clubsuit$ Strong-Field Gravity:
Close to the black hole, spacetime curvature intensifies. This alters the balance between gravitational pull and centrifugal forces, leading to unexpected orbital behaviors.
$\clubsuit$ Orbital Angular Velocity:
Under Newtonian gravity, angular velocity $\Omega$ scales monotonically with radius as:
$\Omega_{\text{Newton}}(r)=\sqrt{\frac{GM}{r^{3}}}\quad\text{(monotonically decreasing with }r\text{)}.$
This implies that as radius decreases, orbital velocity increases uniformly and vice versa.
[…]
Near a rapidly rotating Kerr black hole, the angular velocity is given by 13 :
$\Omega_{\text{Kerr}}(r)=\frac{\sqrt{GM}\left(r^{3/2}+a\sqrt{GM}\right)}{r^{2}\left(r^{3/2}-3\sqrt{GM}r^{1/2}+2a\sqrt{GM}\right)}.$
The numerator includes a spin-enhanced term $a\sqrt{GM}$ that boosts $\Omega(r)$ in prograde orbits. However, the denominator contains relativistic correction terms $(-3\sqrt{GM}r^{1/2})$ linked to intense spacetime curvature, which dampens the angular velocity at small radii.
In summary, for the Radial Evolution of $\Omega(r)$ we have:
- Phase 1 (Large $r$):
Keplerian-like increase in $\Omega$ as $r$ decreases ($d\Omega/dr>0$).
-Phase 2 (Intermediate $r$):
@@
Frame-dragging dominates → $\Omega$ rises super-Keplerianly.
- Phase 3 (Critical Radius $r_{\text{peak}}$):
Spacetime curvature overtakes → $\Omega$ decreases as $r$ decreases ($d\Omega/dr<0$):
$\frac{d\Omega}{dr}>0\quad\xrightarrow{r=r_{\text{peak}}}\quad\frac{d\Omega}{dr}<0.$
@@
This unexpected reversal — the peak and drop in orbital angular velocity at close proximity to the event horizon — which is in sharp contrast with Newtonian expectations is the hallmark of the Aschenbach effect and in essence, arises from the interplay between intense gravitational curvature and spin-induced frame dragging.
Building upon our earlier discussion, we have sawed that the Aschenbach effect—manifested as a non-monotonic profile in the orbital angular velocity—is a direct consequence of frame dragging coupled with strong-field gravity in Kerr black holes. This naturally leads to a fundamental question: Is this behavior unique to rotating black holes, or can static black holes—devoid of intrinsic spin and hence lacking frame dragging—reproduce similar non-monotonic features, albeit through a different mechanism?
@@
In response to this question, we argued in our previous work that while static black holes lack rotational effects, they retain strong gravitational curvature. Consequently, the emergence of a non-monotonic angular velocity profile depends on identifying a surrogate mechanism that replicates the dynamical role of frame dragging. We proposed that this surrogate is the appearance of a stable gravitational potential minimum—either local or global—outside the event horizon, typically associated with a stable photon sphere 14 .
Historically, such “minimum-maximum” potential pairs have been well documented in horizonless ultra-compact gravitational configurations, such that it can be stated with reasonable accuracy that most ultracompact gravitational models in the form of naked singularities exhibit stable photon spheres15 ; 16 ; 17 . However, when the Weak Cosmic Censorship Conjecture (WCCC) is enforced, introducing an event horizon converts these models into legitimate black holes. In most cases, the previously accessible minimum gets hidden behind the horizon, resulting in purely monotonic velocity profiles. Nevertheless, there are models wherein despite the presence of an event horizon, a stable potential minimum persists in the exterior geometry. In such cases, the orbital angular velocity profile exhibits pronounced non-monotonic behavior, centered around this minimum 18 ; 19 . One such example, demonstrated in our earlier research14 , involves a black hole model constructed within a hybrid framework combining Gauss–Bonnet corrections and Massive Gravity. Encouraged by similar findings in another Massive Gravity formulation 19 , we hypothesized that a broad class of black hole geometries within Massive Gravity may serve as viable candidates for exhibiting Aschenbach-like behavior.
To test this hypothesis, we consider three distinct black hole models in Massive Gravity and investigate the conditions under which this phenomenon may arise. Given that the existence and characteristics of potential extrema (such as photon spheres) are highly sensitive to the model’s parametric configuration, our approach entails:
An initial classification of generic black hole traits (existence of event horizons, photon spheres, and TCO behavior),
Followed by a targeted analysis of whether the angular velocity profile develops a non-monotonic signature indicative of the Aschenbach-like effect.
It is important to stress that while a change in the angular velocity profile is the common observable across both rotating and static configurations, the underlying mechanisms are fundamentally different:
@@
In rotating (Kerr) black holes, the Aschenbach effect stems from the interplay between frame dragging and strong gravity.
@@
In static black holes, the analogous behavior is induced through the formation of an exterior gravitational potential minimum, absent frame dragging.
Therefore, although the observational manifestations may resemble one another, the governing physics diverge substantially. This distinction in static models leads us to call this phenomenon in static black holes the Aschenbach-like effect instead of the Aschenbach effect.
II Methodology
@@
In this Article, for the analysis of photon spheres, we adopt the topological charge-based and approach, which has been thoroughly detailed in 20 ; 21 and our previous works 15 ; 16 . In the interest of brevity and to avoid redundancy, we do not revisit the foundational concepts here.
@@
Also, in examining TCOs , we rely on the orbital relations and stability criteria formulated in studies 22 ; 23 ; 24 . Given the comprehensive nature of these references, we direct interested readers to those publications for an in-depth treatment of the methodology.
II.1 Topological Photon Sphere
@@
To study the photonic sphere, we first need an effective potential. We use the effective potential that was constructed for the study of the topological method based on the metric coefficients of a spherically symmetric four-dimensional model. we have 20 ; 21 :
[…]
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}}=\frac{1}{\sin\theta}\bigg{(}\frac{f(r)}{h(r)}\bigg{)}^{1/2}.\end{split}$ (2)
@@
Now, to determine the charges and topological behavior of the model, we will first need a vector field using the effective potential $H$. For this purpose, we consider a general two dimensional vector field as $\phi$ which can be decomposed into two components, $\phi^{r}$ and $\phi^{\Theta}$ 20 ; 21 ,
$\phi=(\phi^{r},\phi^{\Theta}),$ (3)
This vector field can be rewrite as $\phi=||\phi||e^{i\Theta}$, where $||\phi||=\sqrt{\phi\cdot\phi}$. Accordingly, the normalized vector can be defined as,
$n^{\iota}=\frac{\phi^{\iota}}{||\phi||},$ (4)
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}H,\\ &\phi^{\Theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}H}{\sqrt{h(r)}}.\end{split}$ (5)
By employing the previously established relations and the definition of topological charge, one can systematically assign a charge to each photon sphere present in the spacetime 15 ; 16 ; 21 . Before drawing broader conclusions from this section, it is important to underscore a key finding from our earlier publications: In configurations where the model exhibits black hole characteristics, typically satisfying the condition of possessing an event horizon, one often finds a single unstable photon sphere located outside the horizon, characterized by a net topological charge of $-1$. From the perspective of the effective potential, this corresponds to the presence of at least one maximum outside the horizon. Conversely, in horizonless configurations (i.e., naked singularities), the geometry commonly features not only an unstable photon sphere with a topological charge of $-1$, but also one or two stable photon spheres with topological charges of +1. These are identified as local minima of the potential structure and represent dynamically stable light rings. These distinctions are of fundamental importance and will be revisited in subsequent discussions, particularly when analyzing the relationship between topological photon sphere and orbital dynamics in strongly curved spacetimes.
II.2 Time-like Circular Orbits (TCOs)
The geodesic behavior in any gravitational model is profoundly shaped by the form of the effective potential, derived from the interplay of fields within the model’s action. These geodesics may either lack recognizable turning points or exhibit them, forming closed trajectories. In the case of null geodesics, such closed paths correspond to photon spheres, whereas in timelike geodesics, they manifest as TCOs. Both types of orbits may be either stable or unstable, depending on the curvature and potential landscape of the spacetime. In this study, we adopt a static, axisymmetric spacetime with an imposed $\mathbb{Z}_{2}$ symmetry under a (1+3) dimensional framework. This symmetry structure ensures that focusing the analysis on the equatorial plane does not compromise the generality of our results. Based on the metric equation—namely Eq. (1)—we consider the following key quantities 22 :

### v3

# Furthermore, can this Aschenbach-like phenomenon in static configurations be considered as an observable signature in the dynamics of general relativity, similar to the original Aschenbach effect in rotating spacetimes?

@@ Abstract
The Aschenbach effect is widely regarded as a manifestation of two quintessential relativistic features: frame dragging and extreme spacetime curvature. Traditionally associated with rotating geometries, this non-monotonic behavior in orbital angular velocity challenges Newtonian intuition. In our previous work [11], however, we demonstrated that this velocity irregularity is not exclusive to spinning spacetimes. Specifically, we showed that the presence of a stable minimum in the gravitational potential, corresponding to a stable photon sphere, can reproduce Aschenbach-like behavior in static black holes as well. This observation suggests that, even in the absence of rotational frame dragging, curvature alone—if encoded through appropriate geometric extrema—may be sufficient to induce non-monotonic velocity profiles.
@@
keywords: Aschenbach effect, photon spheres, Time-like Circular Orbits, black holes, Massive Gravity
Can Static Black Holes in Massive Gravity Serve as Candidates for Aschenbach-Like Phenomena?
Mohammad Ali S. Afshar  [^1] Email:   m.a.s.afshar@gmail.com,    Jafar Sadeghi  [^2] Email:   pouriya@ipm.ir,    Tahereh Azizi  [^3] Email:   t.azizi@umz.ac.ir,    A. S. Sefiedgar  [^4] Email:   a.sefiedgar@umz.ac.ir
Department of Physics, Faculty of Basic Sciences,
University of Mazandaran P. O. Box 47416-95447, Babolsar, Iran
School of Physics, Damghan University, P. O. Box 3671641167, Damghan, Iran
Canadian Quantum Research Center, 32 Ave, 204-3002 V1T 2L7, Vernon, BC, Canada
1 Introduction
1.1 The Aschenbach Effect: Significance and Origins
@@
The Aschenbach effect represents a compelling manifestation of general relativity in astrophysical con texts. As a tangible prediction of Einstein’s theory in the strong-field regime, observing this phenomenon would confirm that spacetime near black holes behaves precisely as the field equations predict [1, 2, 3] . the systematic analysis of orbital stability near Kerr black holes gained momentum in the late 20th century. Key figures such as Brandon Carter, and later Marek A. Abramowicz, made substantial contributions to the mathematical formalism underpinning these orbital dynamics [4, 5, 6]. Accordingly, This effect was first identified in 2004 when Aschenbach discovered, through numerical simulations, an anomalous drop in orbital velocity for test particles approaching the event horizon of rapidly rotating Kerr black holes [7]. This non-monotonic behavior directly challenges traditional Keplerian flow models com monly employed in simplified accretion disk theories. Unlike quasi-periodic oscillations (QPOs), which typically require magneto-hydrodynamic (MHD) simulations for accurate modeling [8], the Aschenbach effect is purely kinematic and emerges directly from Kerr geometry. Rigorous inferences from Einstein’s field equations demonstrate that the phenomenon originates from the interplay between gravitational forces and rotational effects encoded in the Kerr metric.
1.2 Fundamental Physical Elements
Innermost Stable Circular Orbit (ISCO):
@@ 1.2 Fundamental Physical Elements
This represents the smallest radius at which a test particle can maintain a stable circular orbit around a black hole. Beyond this threshold, minor perturbations trigger inward plunges. The ISCO effectively delineates the inner edge of accretion disks, with matter interior to this radius undergoing rapid radial infall. In rotating spacetimes, ISCO properties depend significantly on orbital orientation relative to the black hole’s spin [9].
Frame Dragging:
@@
The rotation of a Kerr black hole ”drags” surrounding spacetime, altering the dynamics of nearby matter. This effect, quantified by the spin parameter $a=J/M$ (where $J$ is angular momentum and $M$ is mass), is strongest near the event horizon. For maximal spin ($a/M=1$), frame dragging reaches its extremum; it vanishes completely for non-spinning Schwarzschild black holes ($a=0$) [9].
Strong-Field Gravity and Angular Velocity:
In Newtonian gravity, angular velocity $\Omega$ scales monotonically with radius as $\Omega_{\text{Newton}}(r)=\sqrt{GM/r^{3}}$. However, near rapidly rotating Kerr black holes, the angular velocity follows $\Omega_{\text{Kerr}}(r)=\sqrt{M}/(r^{3/2}+a\sqrt{M})$, revealing a complex radial evolution [10]:
1. At large radii: Keplerian-like increase in $\Omega$ as $r$ decreases ($d\Omega/dr<0$)
@@
2. At intermediate radii: Frame-dragging dominates, causing $\Omega$ to rise super-Keplerianly
3. At a critical radius $r_{\text{peak}}$: Spacetime curvature overtakes rotational effects, causing $\Omega$ to decrease as $r$ decreases ($d\Omega/dr>0$)
@@ 2. At intermediate radii: Frame-dragging dominates, causing $\Omega$ to rise super-Keplerianly
This unexpected reversal—the peak and drop in orbital angular velocity near the event horizon—defines the Aschenbach effect and arises from the interplay between intense gravitational curvature and spin-induced frame dragging. While initially identified in asymptotically flat spacetimes, this phenomenon has also been rigorously investigated in asymptotically anti-de Sitter (AdS) and de Sitter (dS) geometries [7].
1.3 Beyond Rotation: Static Configurations and the Aschenbach-like Effect
A fundamental question:
Is non-monotonic angular velocity exclusive to rotating black holes, or can static black holes—devoid of intrinsic spin and frame dragging—reproduce similar behavior through alternative mechanisms?
@@ 1.3 Beyond Rotation: Static Configurations and the Aschenbach-like Effect
In our previous work [11], we demonstrated that while static black holes lack rotational effects, they retain strong gravitational curvature. We proposed that the emergence of a stable gravitational potential minimum—either local or global—outside the event horizon, typically associated with a stable photon sphere, could serve as a surrogate mechanism replicating dynamical aspects of frame dragging.
Although historically, such ”minimum-maximum” potential structures have been well-documented in horizonless ultra-compact objects, with most naked singularity models exhibiting stable photon spheres [12, 13, 14]. However, when the Weak Cosmic Censorship Conjecture (WCCC) is enforced, introducing an event horizon typically hides these minima behind the horizon, resulting in purely monotonic velocity profiles. Nevertheless, certain models maintain exterior potential minima despite the presence of an event horizon, yielding non-monotonic angular velocity profiles centered around these extrema [15, 16].
1.4 Massive Gravity as a Testing Ground
Encouraged by findings in a hybrid Gauss-Bonnet-Massive Gravity black hole model [11] and similar results in another Massive Gravity formulation [16], we hypothesize that black hole geometries within Massive Gravity theories may inherently support Aschenbach-like phenomena. To test this hypothesis, we investigate three distinct black hole models in Massive Gravity, analyzing the conditions under which non-monotonic angular velocity profiles emerge.
@@ 1.4 Massive Gravity as a Testing Ground
Finally, It is crucial to emphasize that while the observational manifestations may resemble one another, the underlying physics differs substantially: - In rotating (Kerr) black holes, the Aschenbach effect stems from frame dragging coupled with strong gravity
@@
- In static black holes, analogous behavior is induced through exterior gravitational potential minima, absent frame dragging
This fundamental distinction justifies our terminological precision: we refer to the phenomenon in static configurations as the ”Aschenbach-like effect” rather than the Aschenbach effect proper. Our investigation seeks to determine whether this Aschenbach-like phenomenon in static configurations can serve as an observable signature of general relativistic dynamics, similar to the original Aschenbach effect in rotating spacetimes.
2 Methodology
@@
In this Article, for the analysis of photon spheres, we adopt the topological charge-based and approach, which has been thoroughly detailed in [17, 18] and our previous works [12, 13]. In the interest of brevity and to avoid redundancy, we do not revisit the foundational concepts here.
@@
Also, in examining TCOs , we rely on the orbital relations and stability criteria formulated in studies [19, 20, 21]. Given the comprehensive nature of these references, we direct interested readers to those publications for an in-depth treatment of the methodology.
2.1 Topological Photon Sphere
@@
To study the photonic sphere, we first need an effective potential. We use the effective potential that was constructed for the study of the topological method based on the metric coefficients of a spherically symmetric four-dimensional model. We have [17, 18]:
$\mathit{ds}^{2}=-\mathit{dt}^{2}f\!\left(r\right)+\frac{\mathit{dr}^{2}}{g\!\left(r\right)}+\left(d\theta^{2}+d\varphi^{2}\sin\!\left(\theta\right)^{2}\right)h\!\left(r\right)=\mathit{dr}^{2}g_{\mathit{rr}}-\mathit{dt}^{2}g_{\mathit{tt}}+d\theta^{2}g_{\theta\theta}+d\varphi^{2}g_{\varphi\varphi},$ (2.1)
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}}=\frac{1}{\sin\theta}\bigg(\frac{f(r)}{h(r)}\bigg)^{1/2}.\end{split}$ (2.2)
@@
Now, to determine the charges and topological behavior of the model, we will first need a vector field using the effective potential $H$. For this purpose, we consider a general two dimensional vector field as $\phi$ which can be decomposed into two components, $\phi^{r}$ and $\phi^{\Theta}$ [17, 18],
$\phi=(\phi^{r},\phi^{\Theta}),$ (2.3)
@@
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}H,\\ &\phi^{\Theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}H}{\sqrt{h(r)}}.\end{split}$ (2.5)
By employing the previously established relations and the definition of topological charge, one can systematically assign a charge to each photon sphere present in the spacetime [12, 13, 18]. Before drawing broader conclusions from this section, it is important to underscore a key finding from our earlier publications: In configurations where the model exhibits black hole characteristics, typically satisfying the condition of possessing an event horizon, one often finds a single unstable photon sphere located outside the horizon, characterized by a net topological charge of 1-1. From the perspective of the effective potential, this corresponds to the presence of at least one maximum outside the horizon. Conversely, in horizonless configurations (i.e., naked singularities), the geometry commonly features not only an unstable photon sphere with a topological charge of 1-1, but also one or two stable photon spheres with topological charges of +1. These are identified as local minima of the potential structure and represent dynamically stable light rings. These distinctions are of fundamental importance and will be revisited in subsequent discussions, particularly when analyzing the relationship between topological photon sphere and orbital dynamics in strongly curved spacetimes.
2.2 Time-like Circular Orbits (TCOs)
The geodesic behavior in any gravitational model is profoundly shaped by the form of the effective potential, derived from the interplay of fields within the model’s action. These geodesics may either lack recognizable turning points or exhibit them, forming closed trajectories. In the case of null geodesics, such closed paths correspond to photon spheres, whereas in timelike geodesics, they manifest as TCOs. Both types of orbits may be either stable or unstable, depending on the curvature and potential landscape of the spacetime. In this study, we adopt a static, axisymmetric spacetime with an imposed $\mathbb{Z}_{2}$ symmetry under a (1+3) dimensional framework. This symmetry structure ensures that focusing the analysis on the equatorial plane does not compromise the generality of our results. Based on the metric equation—namely Eq. (2.1)—we consider the following key quantities [19]:
$A=g_{\varphi\varphi}E^{2}+g_{\mathit{tt}}L^{2},$ (2.6)
$B=-g_{\varphi\varphi}g_{\mathit{tt}},$ (2.7)